OpenStudy (hkmiu):

Rita has deposited $1000 in a bank account. At the end of each year, the bank posts interest to her account in the amount of 5% of the balance, but then takes out a $10 annual fee. would I write it as... bn = 1.05bn - 10 or bn = 1.05b(n-1) - 10 the n is subscript

OpenStudy (hkmiu):

\[b _{n} = 1.05b _{n-1} - 10 \] or \[b _{n} = 1.05b _{n} - 10 \]

ganeshie8 (ganeshie8):

lets label them, \(b_n = \) end of year balance \(b_{n-1} = \) start of year balance

ganeshie8 (ganeshie8):

then, \(\large b_n = 1.05b_{n-1} - 10\) is \(\large \color{red}{\checkmark}\)

OpenStudy (hkmiu):

Whooooo, so I got it write. I'm double checking :P

ganeshie8 (ganeshie8):

also notice that, here initial deposit amount \(\large b_0 = 1000\)

ganeshie8 (ganeshie8):

yes ! good work :)

OpenStudy (hkmiu):

right***********

ganeshie8 (ganeshie8):

using ur formula, after 1 year, the balance wud be, \(\large b_1 = 1.05b_{1-1} - 10\) \(\large b_1 = 1.05b_{0} - 10\) \(\large b_1 = 1.05 \times 1000- 10\) \(\large b_1 = 1040\)

ganeshie8 (ganeshie8):

which seems correct, so your formula is correct too :)

ganeshie8 (ganeshie8):

we can also write the formula like this :- \(\large b_1 = 1000 \) \(\large b_{n+1} = 1.05b_n - 10\)

ganeshie8 (ganeshie8):

i like the above one better, cuz i like the sequence to start wid \(\large b_1\) but your sequence starts wid \(\large b_0\)

OpenStudy (hkmiu):

It actually starts with Bn !

OpenStudy (hkmiu):

Thanks for your help today! I got a 100 on my last few quizzes :). @ganeshie8

ganeshie8 (ganeshie8):

wow ok then you're right :) thats amazing congrats !!!!!!

OpenStudy (hkmiu):

its 4AM, I'm going to sleep. Thanks again :).

ganeshie8 (ganeshie8):

omg ! i bet you deserve some good sleep ! goonight, have horrible math dreams :)

OpenStudy (hkmiu):

Hahahahaa nice oneeee :))))). Bye bye!